Find the value of tan 33° by Lagrange’s formula if

tan 30° = 0.5774, tan 32° = 0.6249,

tan 35° = 0.7002, tan 38° = 0.7813.


1
Expert's answer
2021-01-03T15:56:34-0500

Here the intervals are unequal.


"\begin{matrix}
 x_0=30 & y_0=0.5774 \\
\\
 x_1=32 & y_1=0.6249 \\
\\
 x_2=35 & y_2=0.7002 \\
\\
 x_3=38 & y_3=0.7813 \\
\end{matrix}"

By Lagrange’s interpolation formula we have

"y=f(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)}\times y_3"



Put "x=33"

"f(33)=\dfrac{(33-32)(33-35)(33-38)}{(30-32)(30-35)(30-38)}\times 0.5774"

"+\dfrac{(33-30)(33-35)(33-38)}{(32-30)(32-35)(32-38)}\times 0.6249"


"+\dfrac{(33-30)(33-32)(33-38)}{(35-30)(35-32)(35-38)}\times 0.7002"

"+\dfrac{(33-30)(33-32)(33-35)}{(38-30)(38-32)(38-35)}\times 0.7813"

"=0.6494"

"\tan33\degree=0.6494"